# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ 3.2. The addition of a finite number of point masses
The addition of point masses follows a simple procedure, see, for example, [14]. Diagonal and triangular rational modifications are made to the original RH problem, with the second applying only in a neighborhood of each of the point masses, converting a residue condition (1.6) into a rational jump condition. The effect of this is a rational modification of the densities on $[a_{j},b_{j}]$ for every $j$, a change in the asymptotics at infinity and the introduction of a jump condition on a small curve encircling the point masses. Once the $g$-function is introduced, this jump condition becomes exponentially close to the identity. We sketch this procedure here but refer the reader to [14] for more detail.
@@
$\check{\mathbf{Y}}_{n}(z)=\hat{\mathbf{Y}}_{n}(z)\begin{cases}\begin{bmatrix}1&0\\ -\frac{\tilde{w}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
where $\tilde{w}_{j}$ is defined as
(3.3) $\tilde{w}_{j}:=\frac{2\pi\mathrm{i}}{w_{j}}\prod_{k\neq j}(c_{j}-c_{k})^{-2}.$
@@
    $\check{\mathbf{Y}}_{n}^{+}(z)=\check{\mathbf{Y}}_{n}^{-}(z)\begin{bmatrix}1&0\\ -\frac{\tilde{w}_{j}}{z-c_{j}}&1\end{bmatrix}$
@@
$\mathbf{H}^{\natural}(x,y)$ =$=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm},\vspace{4pt}\\ \begin{bmatrix}1&0\\ \frac{\tilde{w}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\rho$ is replaced with (3.4) and $n$ is replaced with $n-P$. Similarly, Theorem 3.1 holds if one makes the new definition
@@
$\mathbf{H}^{\natural}(x,y)$ =$=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega^{\prime}_{j,\pm},\vspace{4pt}\\ \begin{bmatrix}1&0\\ \frac{\tilde{w}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\rho$ is replaced with (3.4) and $n$ is replaced with $n-P$.
@@ 4. Asymptotics of recurrence coefficients and estimating optimal errors
To compare our error estimates to the true errors, we consider a single interval $[a_{1},b_{1}]=[-1,1]$, $g=0$. We use $\alpha_{1}=\beta_{1}=1/2$,
@@
$\int_{0}^{1}p(x)x^{k+\alpha}\sqrt{1-x}\mathop{}\!\mathrm{d}x$ $=\sum_{j=1}^{n}p(y_{j})v_{j},$
@@
when $f$ and $g$ are polynomials of degree less than or equal to $n-1$. This implies the recurrence coefficients can be generated exactly (up to roundoff). The ideal algorithm is the RKPW method as outlined in [21] which has $O(n^{2})$ complexity. In Figures 4 and 5 we provide evidence that error term in the asymptotics of the recurrence coefficients is $O(n^{-\gamma-1})$ whereas our method using $q=\infty$ predicts $O(n^{-\gamma})$.
@@ Appendix A The $g$-function
In this section we define what is commonly called the external Green’s function with pole at infinity [35] and determine some essential properties.
Specifically, we find a function $\mathfrak{g}$ that satisfies:
@@
    $\mathfrak{g}_{+}^{\prime}(z),\mathfrak{g}_{-}^{\prime}(z)\in\mathrm{i}\mathbb{R}$ on $[a_{j},b_{j}]$.
@@
    $\int_{b_{j}}^{a_{j+1}}\mathfrak{g}(z)\mathop{}\!\mathrm{d}z=0$, $j=1,2,\ldots,g$.
@@
    $\mathfrak{g}^{+}(z)-\mathfrak{g}^{-}(z)=:\Delta_{j}\in\mathrm{i}\mathbb{R},$ is constant on $[b_{j},a_{j+1}]$, and
@@
There exists $\epsilon>0$ and $\gamma>0$ such that if $a_{j}-\epsilon\leq x\leq b_{j}+\epsilon$ and $|y|\leq\epsilon$ then
$\,\mathrm{e}^{\operatorname{Re}\mathfrak{g}(z)}>|\phi(M(z;a_{j},b_{j}))|^{\gamma}.$
@@ Appendix B Extension operator
$m(z;c)=\sum_{\ell=0}^{\kappa}\frac{f_{j}^{(\ell)}(c)}{\ell!}z^{\ell},\quad c=a,b.$
For $20<\theta_{1}<\theta_{2}<\pi/$ 2, construct an infinitely smooth function that for $\theta\in[-\pi/2,\pi/2]$ satisfies
@@
Then extend $\tilde{\beta}(\theta)$ to the entirety of interval $[-\pi,\pi]$ using an even reflection about $\theta=\pi/2$. Then, one obtains a function on the unit circle in the obvious way:
@@
(B.3) $\begin{split}m^{\sharp}(x,y)&=\left[\tilde{b}(x)m(x,y)\right.\\ &+(1-\tilde{b}(x))(1-\beta(x-a,y))m(x,y)+\beta(x-a,y)(1-\tilde{b}(x))m(z;a)\\ &\left.+(1-\tilde{b}(x))(1-\beta(x-b,y))m(x,y)+\beta(x-b,y)(1-\tilde{b}(x))m(z;b)\right](1-b_{\tau}(|y|;0)).\end{split}$
@@
    $m^{\sharp}(x,y)$ $=m(z;a),\quad\theta_{2}<|\arg(z-a)|<\pi,$
    $m^{\sharp}(x,y)$ $=m(z;b),\quad\theta_{2}<|\arg(b-z)|<\pi,$
@@
  4. (4)
    for $a\leq x\leq b$ and $|y|\leq\tau$, ¯$\|\overline{\partial}m^{\sharp}(\cdot,y)\|_{L^{\infty}([a,b])}\leq C_{\kappa}\|m\|_{C^{\kappa,1}([a,b])}|y|^{\kappa}$, for some $C_{\kappa}>0$, and
@@ B.1. Verification of the properties
$\left|\sum_{j=0}^{\kappa}\frac{f^{(j)}(x)}{j!}(\mathrm{i}y)^{j}-\sum_{j=0}^{\kappa}\frac{f^{(j)}(0)}{j!}(x+\mathrm{i}y)^{j}\right|\leq C_{\kappa}\|f\|_{C^{\kappa,1}(I)}\sum_{j=0}^{\kappa}|x|^{j}|y|^{\kappa+1-j}$
@@
where $|R_{j}(x)|\leq M_{j}\sup_{\xi\in I}|f^{(\kappa-j+1)}(\xi)||x|^{\kappa-j+1}$. Then, the binomial theorem gives
@@ B.1.1. Property (1)
For $a+2\tau\leq x\leq b-2\tau$ we have
$m^{\sharp}(x,0)=m(x,0)=m(x),$
because $\tilde{b}(x)=1$. Then for $a\leq x\leq a+2\tau$
$m^{\sharp}(x,0)=\tilde{b}(x)m(x,0)+(1-\tilde{b}(x))m(x,0)$
A similar calculation follows near $x=b$.
B.1.2. Property (2)
By definition, for $|z-a|<\tau$ and $\theta_{2}<|\arg(z-a)|<\pi$ we find that $m^{\sharp}(x,y)=m(z;a)$. The same calculation follows near $x=b$.
B.1.3. Property (3)
@@ B.1.4. Property (4)
So, we immediately have the property we desire for $z$ such that $a+2\tau\leq x\leq b-2\tau$, when $\tilde{b}(x)=1$. Then for $|z-a|\leq 2\tau$ and $-\theta_{1}<\arg(z-a)<\theta_{1}$
@@
Since ¯$\overline{\partial}m(z;a)=0$ we find
$2\overline{\partial}m^{\sharp}(x,y)$ $=-2\beta(x-a,y)\overline{\partial}m(x,y)+2\overline{\partial}\beta(x-a,y)\left[m(z;a)-m(x,y)\right].$
@@
It remains to bound ¯$\overline{\partial}\beta$.
@@
The establishes Property (4), with a $\tau$-independent constant for $|y|\leq\tau/2$.
@@ Appendix C Construction and estimation of local parametrices
(C.1) $\mathbf{P}_{\operatorname{Bessel}}(\xi;\alpha)=\begin{cases}\begin{bmatrix}\operatorname{I}_{\alpha}(2\xi^{1/2})&\frac{\mathrm{i}}{\pi}\operatorname{K}_{\alpha}(2\xi^{1/2})\\ 2\pi\mathrm{i}\xi^{1/2}\operatorname{I}_{\alpha}^{\prime}(2\xi^{1/2})&-2\xi^{1/2}\operatorname{K}_{\alpha}^{\prime}(2\xi^{1/2})\end{bmatrix}&\arg\xi\in(-\sigma,\sigma),\vspace{4pt}\\ \begin{bmatrix}\frac{1}{2}\operatorname{H}_{\alpha}^{(1)}(2(-\xi)^{1/2})&\frac{1}{2}\operatorname{H}_{\alpha}^{(2)}(2(-\xi)^{1/2})\\ \pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(1)}}^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(2)}}^{\prime}(2(-\xi)^{1/2})\end{bmatrix}\,\mathrm{e}^{\frac{1}{2}\alpha\pi\mathrm{i}{\boldsymbol{\sigma}_{3}}}&\arg\xi(\sigma,\pi),\vspace{4pt}\\ \begin{bmatrix}\frac{1}{2}\operatorname{H}_{\alpha}^{(2)}(2(-\xi)^{1/2})&-\frac{1}{2}\operatorname{H}_{\alpha}^{(1)}(2(-\xi)^{1/2})\\ -\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(2)}}^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(1)}}^{\prime}(2(-\xi)^{1/2})\end{bmatrix}\,\mathrm{e}^{-\frac{1}{2}\alpha\pi\mathrm{i}{\boldsymbol{\sigma}_{3}}}&\arg\xi\in(-\pi,-\sigma).\end{cases}$
@@
$\mathbf{J}_{\operatorname{Bessel}}(\xi;\alpha)$ =$=\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{\alpha\pi\mathrm{i}}&1\end{bmatrix}&\xi\in\Gamma_{1},\vspace{4pt}\\ \ \begin{bmatrix}0&1\\ -1&0\end{bmatrix}&\xi\in\Gamma_{2},\vspace{4pt}\\ \ \begin{bmatrix}1&0\\ \,\mathrm{e}^{-\alpha\pi\mathrm{i}}&1\end{bmatrix}&\xi\in\Gamma_{3},\end{cases}$
where contours are oriented in the direction of increasing modulus. The asymptotics at the origin are given by the following three cases. If $\alpha>0$
$\mathbf{P}_{\operatorname{Bessel}}(\xi;\alpha)$ =$=\begin{cases}\mathrm{O}\begin{bmatrix}|\xi|^{\alpha/2}&|\xi|^{-\alpha/2}\\ |\xi|^{\alpha/2}&|\xi|^{-\alpha/2}\end{bmatrix}&|\arg\xi|<2\pi/3,\\ \\ \mathrm{O}\begin{bmatrix}|\xi|^{-\alpha/2}&|\xi|^{-\alpha/2}\\ |\xi|^{-\alpha/2}&|\xi|^{-\alpha/2}\end{bmatrix}&2\pi/3<|\arg\xi|<\pi,\end{cases}\text{ as }\xi\to 0.$
@@
We also remark that $\det\mathbf{P}_{\operatorname{Bessel}}=1$ (see [28, 10.28.2]) so that asymptotics for the inverse of $\mathbf{P}_{\operatorname{Bessel}}$ can be easily inferred. We also have
@@
×$\times(\mathbf{I}+\mathrm{O}(n^{-1}|\xi|^{-1/2})),~{}~{}n\to\infty.$
@@
(C.2) $\begin{split}\mathbf{P}_{\operatorname{Bessel}}(n^{2}\xi;\alpha)&=(\pi n)^{-\frac{1}{2}{\boldsymbol{\sigma}_{3}}}\xi^{-\frac{1}{4}{\boldsymbol{\sigma}_{3}}}\mathbf{E}_{\operatorname{Bessel}}(n^{2}\xi)\,\mathrm{e}^{2n\xi^{1/2}{\boldsymbol{\sigma}_{3}}},\\ \mathbf{E}_{\operatorname{Bessel}}(n^{2}\xi;\alpha)&=\mathbf{E}+\mathrm{O}(n^{-1}|\xi|^{-1/2}),\quad\mathbf{E}=\begin{bmatrix}\frac{1}{2}&\frac{\mathrm{i}}{2}\\ \mathrm{i}&1\end{bmatrix}.\end{split}$
@@
$W(z;a_{j})$ $=(z-b_{j})_{\rightarrow}^{\alpha_{j}/2}(z-a_{j})_{\rightarrow}^{\beta_{j}/2}f(z;a_{j})^{-1/2},$
@@
$\mathfrak{g}(z)=\int_{a_{1}}^{b_{j}}\mathfrak{g}_{+}^{\prime}(z)\mathop{}\!\mathrm{d}z+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z^{\prime})\mathop{}\!\mathrm{d}z^{\prime}=\sum_{\ell=1}^{j}\int_{a_{\ell}}^{b_{\ell}}\mathfrak{g}_{+}^{\prime}(z)\mathop{}\!\mathrm{d}z+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z)\mathop{}\!\mathrm{d}z=\frac{\Delta_{j}}{2}+\int_{b_{j}}^{z}\mathfrak{g}_{+}^{\prime}(z).$
@@
$\mathfrak{g}(z)=-\frac{\Delta_{j}}{2}+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z).$
@@
×$\times\,\mathrm{e}^{-(\operatorname{sign}\operatorname{Im}z)\frac{n}{2}\Delta_{j-1}{\boldsymbol{\sigma}_{3}}}W(z;a_{j})^{-{\boldsymbol{\sigma}_{3}}}.$
@@
×$\times\,\mathrm{e}^{-(\operatorname{sign}\operatorname{Im}z)\frac{n}{2}\Delta_{j}{\boldsymbol{\sigma}_{3}}}W(z;b_{j})^{-{\boldsymbol{\sigma}_{3}}}.$

### v4

@@ 3.2. The addition of a finite number of point masses
The addition of point masses follows a simple procedure, see, for example, [17]. Diagonal and triangular rational modifications are made to the original RH problem, with the second applying only in a neighborhood of each of the point masses, converting a residue condition (1.6) into a rational jump condition. The effect of this is a rational modification of the densities on $a_{j},b_{j}]$ for every $j$, a change in the asymptotics at infinity and the introduction of a jump condition on a small curve encircling the point masses. Once the $g$-function is introduced, this jump condition becomes exponentially close to the identity. We sketch this procedure here but refer the reader to [17] for more detail.
@@
$\check{\mathbf{Y}}_{n}(z)=\hat{\mathbf{Y}}_{n}(z)\begin{cases}\begin{bmatrix}1&0\\ -\frac{\tilde{r}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
where $\tilde{r}_{j}$ is defined as
(3.3) $\tilde{r}_{j}:=\frac{2\pi\mathrm{i}}{{r}_{j}}\prod_{k\neq j}(c_{j}-c_{k})^{-2}.$
@@
    $\check{\mathbf{Y}}_{n}^{+}(z)=\check{\mathbf{Y}}_{n}^{-}(z)\begin{bmatrix}1&0\\ -\frac{\tilde{r}_{j}}{z-c_{j}}&1\end{bmatrix}$
@@
$\mathbf{H}^{\natural}(x,y)$ $=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm},\vskip 4.0pt\\ \begin{bmatrix}1&0\\ \frac{\tilde{r}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\rho$ is replaced with (3.4) and $n$ is replaced with nPn-P. Similarly, Theorem 3.1 holds if one makes the new definition
@@
$\mathbf{H}^{\natural}(x,y)$ $=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega^{\prime}_{j,\pm},\vskip 4.0pt\\ \begin{bmatrix}1&0\\ \frac{\tilde{r}_{j}}{z-c_{j}}&1\end{bmatrix}&z\in\Xi_{j},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\rho$ is replaced with (3.4) and $n$ is replaced with nPn-P.
@@ 4. Asymptotics of recurrence coefficients and estimating optimal errors
To compare our error estimates to the true errors, we consider a single interval $a_{1},b_{1}]=[-1,1]$, $g=0$. We use $\alpha_{1}=\beta_{1}=1/2$,
@@
$\int_{0}^{1}p(x)x^{\gamma}\sqrt{1-x^{2}}\mathop{}\!\mathrm{d}x$ $=\sum_{j=1}^{n}p(y_{j})v_{j},$
@@
when $f$ and $g$ are polynomials of degree less than or equal to n1n-1. This implies the recurrence coefficients can be generated exactly (up to roundoff). The ideal algorithm is the RKPW method as outlined in [24] which has $O(n^{2})$ complexity. In Figures 4 and 5 we provide evidence that error term in the asymptotics of the recurrence coefficients is $O(n^{-\gamma-1})$ whereas our method, and that of Yattselev [40], using $q=\infty$ predicts $O(n^{-\gamma}\log n)$.
@@ Appendix A The $g$-function
In this section we define what is commonly called the external Green’s function with pole at infinity [38] and determine some essential properties. This construction can be also be found in [2]. Specifically, we find a function $\mathfrak{g}$ that satisfies:
@@
    $\mathfrak{g}_{+}^{\prime}(z),\mathfrak{g}_{-}^{\prime}(z)\in\mathrm{i}\mathbb{R}$ on $a_{j},b_{j}]$.
@@
    $\int_{b_{j}}^{a_{j+1}}{\mathfrak{g}^{\prime}}(z)\mathop{}\!\mathrm{d}z=0$, $j=1,2,\ldots,g$.
@@
    $\mathfrak{g}^{+}(z)-\mathfrak{g}^{-}(z)=:\Delta_{j}\in\mathrm{i}\mathbb{R},$ is constant on $b_{j},a_{j+1}]$, and
@@
There exists $\epsilon>0$ and $\gamma,\gamma^{\prime}>0$ such that if $a_{j}-\epsilon\leq x\leq b_{j}+\epsilon$ and $|y|\leq\epsilon$ then
${\,\mathrm{e}^{\operatorname{Re}\mathfrak{g}(z)}\geq|\phi(M(z;a_{j},b_{j}))|^{\gamma}\quad\text{and}\quad\,\mathrm{e}^{\operatorname{Re}\mathfrak{g}(z)}\leq|\phi(M(z;a_{j},b_{j}))|^{\gamma^{\prime}}.}$
@@ Appendix B Extension operator
$m(z;c)=\sum_{\ell=0}^{\kappa}\frac{{m}^{(\ell)}(c)}{\ell!}z^{\ell},\quad c=a,b.$
For $0<\theta_{1}<\theta_{2}<\pi/2$, construct an infinitely smooth function that for $\theta\in[-\pi/2,\pi/2]$ satisfies
@@
Then extend $\tilde{\beta}(\theta)$ to the entirety of interval $\pi,\pi]$ using an even reflection about $\theta=\pi/2$. Then, one obtains a function on the unit circle in the obvious way:
@@
(B.3) $\begin{split}m^{\sharp}(x,y)&=(1-b_{\tau}(|y|;0))\left[\tilde{b}(x)m(x,y)\right.\\ &+(1-b_{\tau}(x;a)(1-\beta(x-a,y))m(x,y)+\beta(x-a,y)(1-b_{\tau}(x;a))m(z;a)\\ &\left.+(1-b_{\tau}(x;b))(1-\beta(x-b,y))m(x,y)+\beta(x-b,y)(1-b_{\tau}(x;b))m(z;b)\right].\end{split}$
@@
    $m^{\sharp}(x,y)$ $=m(z;a),\quad\theta_{2}<|\arg(z-a)|<{\pi/2},$
    $m^{\sharp}(x,y)$ $=m(z;b),\quad\theta_{2}<|\arg(b-z)|<{\pi/2},$
@@
    for $a\leq x\leq b$ and $|y|\leq\tau$, $\|\overline{\partial}m^{\sharp}(\cdot,y)\|_{L^{\infty}([a,b])}\leq C_{\kappa}\|m\|_{C^{\kappa,1}([a,b])}|y|^{\kappa}$, for some $C_{\kappa}>0$, and
@@ B.1. Verification of the properties
$\left|\sum_{j=0}^{\kappa}\frac{f^{(j)}(x)}{j!}(\mathrm{i}y)^{j}-\sum_{j=0}^{\kappa}\frac{f^{(j)}(0)}{j!}(x+\mathrm{i}y)^{j}\right|\leq C\|f\|_{C^{\kappa,1}(I)}{\sum_{j=1}^{\kappa+1}}|x|^{j}|y|^{\kappa+1-j}$
@@
where $|R_{j}(x)|\leq M_{j}\sup_{\xi\in I}|f^{({\kappa+1})}(\xi)||x|^{\kappa-j+1}$. Then, the binomial theorem gives
@@ B.1.1. Property (1)
B.1.2. Property (2)
For $a+2\tau\leq x\leq b-2\tau$ we have $b_{\tau}(x;a)=b_{\tau}(x;b)=1$ giving
$m^{\sharp}(x,0)=m(x,0)=m(x).$
Then for $a\leq x\leq a+2\tau$, we have $b_{\tau}(x;b)=1$ and
$m^{\sharp}(x,0)={b_{\tau}(x;a)m(x,0)+(1-b_{\tau}(x;a))m(x,0)},$
A similar calculation follows near $x=b$.
B.1.3. Property (3)
By definition, for $|z-a|<\tau$ and $\theta_{2}<|\arg(z-a)|<{\pi/2}$ we find that $m^{\sharp}(x,y)=m(z;a)$. The same calculation follows near $x=b$.
@@ B.1.4. Property (4)
So, we immediately have the property we desire for $z$ such that $a+2\tau\leq x\leq b-2\tau$, when $\tilde{b}(x)=1$. Then for $|z-a|\leq 2\tau$ and $\theta_{1}<\arg(z-a)<\theta_{1}$
@@
Since $\overline{\partial}m(z;a)=0$ we find
$2\overline{\partial}m^{\sharp}(x,y)$ $={(2-2\beta(x-a,y))}\overline{\partial}m(x,y)+2\overline{\partial}\beta(x-a,y)\left[m(z;a)-m(x,y)\right].$
@@
It remains to bound $\overline{\partial}\beta$.
@@
This establishes Property (4), with a $\tau$-independent constant for $|y|\leq\tau/2$.
@@ Appendix C Construction and estimation of local parametrices
(C.1) $\mathbf{P}_{\operatorname{Bessel}}(\xi;\alpha)=\begin{cases}\begin{bmatrix}\operatorname{I}_{\alpha}(2\xi^{1/2})&\frac{\mathrm{i}}{\pi}\operatorname{K}_{\alpha}(2\xi^{1/2})\\ 2\pi\mathrm{i}\xi^{1/2}\operatorname{I}_{\alpha}^{\prime}(2\xi^{1/2})&-2\xi^{1/2}\operatorname{K}_{\alpha}^{\prime}(2\xi^{1/2})\end{bmatrix}&\arg\xi\in(-\sigma,\sigma),\vskip 4.0pt\\ \begin{bmatrix}\frac{1}{2}\operatorname{H}_{\alpha}^{(1)}(2(-\xi)^{1/2})&\frac{1}{2}\operatorname{H}_{\alpha}^{(2)}(2(-\xi)^{1/2})\\ \pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(1)}}^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(2)}}^{\prime}(2(-\xi)^{1/2})\end{bmatrix}\,\mathrm{e}^{\frac{1}{2}\alpha\pi\mathrm{i}{\boldsymbol{\sigma}_{3}}}&\arg\xi(\sigma,\pi),\vskip 4.0pt\\ \begin{bmatrix}\frac{1}{2}\operatorname{H}_{\alpha}^{(2)}(2(-\xi)^{1/2})&-\frac{1}{2}\operatorname{H}_{\alpha}^{(1)}(2(-\xi)^{1/2})\\ -\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(2)}}^{\prime}(2(-\xi)^{1/2})&\pi\xi^{1/2}{\operatorname{H}_{\alpha}^{(1)}}^{\prime}(2(-\xi)^{1/2})\end{bmatrix}\,\mathrm{e}^{-\frac{1}{2}\alpha\pi\mathrm{i}{\boldsymbol{\sigma}_{3}}}&\arg\xi\in(-\pi,-\sigma).\end{cases}$
@@
$\mathbf{J}_{\operatorname{Bessel}}(\xi;\alpha)$ $=\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{\alpha\pi\mathrm{i}}&1\end{bmatrix}&\xi\in\Gamma_{1},\vskip 4.0pt\\ \ \begin{bmatrix}0&1\\ -1&0\end{bmatrix}&\xi\in\Gamma_{2},\vskip 4.0pt\\ \ \begin{bmatrix}1&0\\ \,\mathrm{e}^{-\alpha\pi\mathrm{i}}&1\end{bmatrix}&\xi\in\Gamma_{3},\end{cases}$
where contours are oriented in the direction of decreasing modulus. The asymptotics at the origin are given by the following three cases. If $\alpha>0$
$\mathbf{P}_{\operatorname{Bessel}}(\xi;\alpha)$ $=\begin{cases}\mathrm{O}\begin{bmatrix}|\xi|^{\alpha/2}&|\xi|^{-\alpha/2}\\ |\xi|^{\alpha/2}&|\xi|^{-\alpha/2}\end{bmatrix}&|\arg\xi|<2\pi/3,\\ \\ \mathrm{O}\begin{bmatrix}|\xi|^{-\alpha/2}&|\xi|^{-\alpha/2}\\ |\xi|^{-\alpha/2}&|\xi|^{-\alpha/2}\end{bmatrix}&2\pi/3<|\arg\xi|<\pi,\end{cases}\text{ as }\xi\to 0.$
@@
We also remark that $\det\mathbf{P}_{\operatorname{Bessel}}=1$ (see [31, 10.28.2]) so that asymptotics for the inverse of $\mathbf{P}_{\operatorname{Bessel}}$ can be easily inferred. We also have
@@
$\times(\mathbf{I}+\mathrm{O}(n^{-1}|\xi|^{-1/2})),~~n\to\infty.$
@@
[…]
@@
$W(z;a_{j})$ $={\,\mathrm{e}^{-\mathrm{i}\pi(\alpha_{j}+\beta_{j})/2}}(z-b_{j})_{\rightarrow}^{\alpha_{j}/2}(z-a_{j})_{\rightarrow}^{\beta_{j}/2}f(z;a_{j})^{-1/2},$
@@
$\mathfrak{g}(z)=\int_{a_{1}}^{b_{j}}\mathfrak{g}_{+}^{\prime}(z)\mathop{}\!\mathrm{d}z^{\prime}+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z^{\prime})\mathop{}\!\mathrm{d}z^{\prime}=\sum_{\ell=1}^{j}\int_{a_{\ell}}^{b_{\ell}}\mathfrak{g}_{+}^{\prime}(z)\mathop{}\!\mathrm{d}z^{\prime}+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z)\mathop{}\!\mathrm{d}z^{\prime}=\frac{\Delta_{j}}{2}+\int_{b_{j}}^{z}\mathfrak{g}_{+}^{\prime}(z)\mathop{}\!\mathrm{d}z^{\prime}.$
@@
$\mathfrak{g}(z)=-\frac{\Delta_{j}}{2}+\int_{b_{j}}^{z}\mathfrak{g}^{\prime}(z)\mathop{}\!\mathrm{d}z^{\prime}.$
@@
$\times\,\mathrm{e}^{-(\operatorname{sign}\operatorname{Im}z)\frac{n}{2}\Delta_{j-1}{\boldsymbol{\sigma}_{3}}}W(z;a_{j})^{-{\boldsymbol{\sigma}_{3}}}.$
@@
$\times\,\mathrm{e}^{-(\operatorname{sign}\operatorname{Im}z)\frac{n}{2}\Delta_{j}{\boldsymbol{\sigma}_{3}}}W(z;b_{j})^{-{\boldsymbol{\sigma}_{3}}}.$